OpenStudy (anonymous):

Write the equation of the line that is tangent to the circle (x – 2)2 + (y – 4)2 = 289, at the point (-15, 4). a. y=4 b. x=-15 c. x=4 d. y=-15

jimthompson5910 (jim_thompson5910):

not 100% sure, but I'm guessing you're not familiar with calculus right?

jimthompson5910 (jim_thompson5910):

especially the concept of derivatives

OpenStudy (anonymous):

No. Is that calculus?

jimthompson5910 (jim_thompson5910):

yes, seems odd how they can ask you this question when you should know about derivatives luckily there's a shortcut

jimthompson5910 (jim_thompson5910):

the center is (2,4) the point on the circle is (-15, 4) so the slope of the line through these two points is m = (y2-y1)/(x2-x1) m = (4-4)/(-15-2) m = 0/(-17) m = 0

jimthompson5910 (jim_thompson5910):

so this means that the line connecting the center to this point is a horizontal line

OpenStudy (anonymous):

okay.

jimthompson5910 (jim_thompson5910):

this must mean that the tangent line is a vertical line (because it makes a right angle with the radius at this point)

jimthompson5910 (jim_thompson5910):

the x coordinate of the point (-15, 4) is x = -15 so the equation of the vertical line is x = -15

jimthompson5910 (jim_thompson5910):

if something isn't clicking, drawing it out may help

OpenStudy (anonymous):

Your a life saver. Thank you for all that. And that's a good idea.

jimthompson5910 (jim_thompson5910):

glad to be of help